OpenStudy (anonymous):

express in exponiential form or radical form (sqaure rooted by p ) ^11

mathslover (mathslover):

is this your question? : \[\large{(\sqrt{p})^{11}}\] @sheetalvee

OpenStudy (anonymous):

yes

mathslover (mathslover):

ok so first of all you must know that : \[\large{\sqrt{a}=a^{\frac{1}{2}}}\]

mathslover (mathslover):

so we can write : \[\large{\sqrt{p}=p^{\frac{1}{2}}}\]

mathslover (mathslover):

now we have : \[\large{(p^{\frac{1}{2}})^{11}}\]

mathslover (mathslover):

Now we must also know that : \[\large{(a^b)^c=a^{bc}}\]

mathslover (mathslover):

hence we have answer as : \[\large{p^{\frac{1}{2}*11}}\]

mathslover (mathslover):

that is all just simplify 2 * 11/2 further..... you will get to your correct answer

OpenStudy (anonymous):

p^11/2

OpenStudy (goformit100):



OpenStudy (anonymous):

thanks

mathslover (mathslover):

Very good and good work and welcome